If P=2y-6xy-3x and Q=4-5y, find the value of 2P-3Q
Found 2 solutions by user_dude2008, MathTherapy:
Answer by user_dude2008(1862)


2P-3Q = 2(2y-6xy-3x)-3(4-5y)
      = 4y-12xy-6x-12+15y
      = 19y-12xy-6x-12




Answer by MathTherapy(10551)

P = 2y-6xy-3x
2P = 2(2y-6xy-3x), or 4y – 12xy – 6x
Q=4-5y
3Q = 3(4 – 5y) or 12 - 15y
2P-3Q = 4y – 12xy – 6x – (12 - 15y)
= 4y – 12xy – 6x – 12 + 15y
= - 6x + 19y – 12xy – 12 Answer.